Prove that [(-a,b)] is the additive inverse for [(a,b)] in the field of quotients. NOTE: these are equivalence classes


1
Expert's answer
2021-09-21T11:30:34-0400

Solution:

We know that in the field of quotients "[(a, b)]=\{(c,d):ad=bc\},\ [(0,1)]" is the additive neutral element and "[(a,b)]+[(c,d)]=[(ad+bc,bd)]." Taking into account that "0\cdot 1=b^2\cdot 0," and hence "[(0,b^2)]=[(0,1)]," we conclude that "[(\u2212a, b)]+[(a, b)]=[(-ab+ba,b^2)]=[(-ab+ab,b^2)]=[(0,b^2)]=[(0,1)]," and therefore "[(\u2212a, b)]" is the additive inverse for "[(a, b)]" in the field of quotients.